OpenStudy (fanduekisses):

Does the (x+5) in( (x+5)-(2x-1))/(x+5) cancel out?

OpenStudy (fanduekisses):

Or no since the 2x-1 is beibg subtracted?

OpenStudy (solomonzelman):

You question is: \(\color{#000000}{ \displaystyle \frac{x+5-(2x-1)}{x+5} }\) Right?

OpenStudy (fanduekisses):

Yes

OpenStudy (solomonzelman):

Then, no, it does not cancel out, because (as you said) 2x-1 is being subtracted.

OpenStudy (fanduekisses):

My intuition tells me no b it idk

OpenStudy (fanduekisses):

Oohhh great hehe

OpenStudy (solomonzelman):

You want to simplify it, correct?

OpenStudy (mathmale):

You have 2 terms in the numerator. You cannot cancel one of them and not cancel the other. So, the answer to your question is ... ?

OpenStudy (solomonzelman):

There are multiple things you can do in such case.... it all depends on your query/need.

OpenStudy (solomonzelman):

You can write it as a polynomial with a remainder (essentially preform long/synthetic division). Or, you can add like terms, and write the fraction ....

OpenStudy (fanduekisses):

Yes distribute them combine like terms

OpenStudy (solomonzelman):

Ok, so first you need to distribute/expand -(2x-1)=?

OpenStudy (mathmale):

The question pertains to cancelling: Is it or is it not possible in this situation? No mathematical operations on the numerator will change that fact. Either you can cancel the (x+5) or you cannot cancel the (x+5). That's all this problem asks you to specify.

OpenStudy (solomonzelman):

Oh, that is exactly the question they are asking .... I have incorrectly read this.